two angles are complementary.the sum of the measure of the 1st angle and half the 2nd angle is 79 degrees.what are the measure of the angles.
Answer by mananth(16949)

let one angle be x
the other angle will be 90-x
x+1/2*(90-x)=79
2x+90-x=2*79
x+90= 158
x=158-90
x=68
the other angle = 22